OpenStudy (anonymous):

The length of a rectangular playing field is 6 meters less than four times its width. If 118meters of fencing goes around the field, find the dimensions of the field.

OpenStudy (anonymous):

l = 4*w - 6 and 2*l+2*w=118

OpenStudy (anonymous):

Maybe you can go from there.

OpenStudy (anonymous):

Does w=23?

OpenStudy (anonymous):

And l=29/4??

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

A hint: w equals 13.